# Pasadena, CA 91125, USA

### v1

@@ Abstract
We propose a new fermionic sum formula for the Macdonald index for a class of Argyres-Douglas theories. The formula arises naturally from a three-dimensional topological field theory obtained by a twisted dimensional reduction of the 4d theory. Such a reduction often gives rise to a 3d $\mathcal{N}=2$ abelian Chern-Simons matter theory that is expected to flow to an $\mathcal{N}=4$ superconformal fixed point. After performing a topological twist, we obtain a 3d TFT admitting boundary conditions that support the vertex operator algebra associated with the original 4d theory. In this framework, the Macdonald index appears as a half-index of the 3d gauge theory, with the Macdonald grading determined by a distinguished $U(1)_{A}$ symmetry in the infrared $\mathcal{N}=4$ superconformal algebra. We present a general procedure to identify this $U(1)_{A}$ symmetry and, whenever possible, show that it produces the known prescription for computing the Macdonald index from the refined character of the associated vertex operator algebra. Our construction also gives a hint towards the IR formula for the Macdonald index, from the 4d BPS particles on the Coulomb branch.
@@ 1 Introduction
$I_{S}(q)=\text{Tr }{\mathcal{O}}(q)\ ,$ (2)
where ${\mathcal{O}}(q)$ is the quantum monodromy operator constructed from the BPS spectrum in the Coulomb branch of the 4d theory. This representation provides a deep connection between UV operator algebras and infrared BPS dynamics, suggesting interrelations among wall-crossing phenomena, protected operator algebras, and the underlying modular structure of supersymmetric field theories.
@@
where the trace is taken over the space of boundary local operators, and $R_{H},~R_{C}$ correspond to Cartan generators of the $SU(2)_{H}\times SU(2)_{C}$ R-symmetry of the $\mathcal{N}=4$ algebra. For a large class of Argyres-Douglas theories, the $A$-twisted half-index obtained in this way is identified with the trace of the quantum monodromy operator (2), which in turn provides a physical motivation for the infrared trace formula of the Schur index. See also ArabiArdehali:2024ysy ; Ardehali:2024 for an alternative approach based on studying the high-temperature limit of the superconformal index.
@@
where the trace is over the operators satisfying the relation $\Delta=2R+2j_{2}-r$ and $j_{1}-j_{2}+r=0$. This is a natural refinement to consider in the context of the 4d SCFT/2d VOA correspondence, since it is built from the same protected operator spectrum as the Schur index while keeping track of an additional $U(1)$ charge. From the perspective of the associated VOA, this refinement arises from the existence of a natural filtration structure, as first discussed in Song:2016yfd . See also Agarwal:2018zqi ; Bonetti:2018fqz ; Beem:2019tfp ; Xie:2019zlb ; Agarwal:2021oyl for related developement.
The 3d $\mathcal{N}=4$ SCFT obtained from the twisted circle reduction has an $\mathcal{N}=2$ superconformal subalgebra where the $U(1)_{R}$ symmetry embeds into $SU(2)_{H}\times SU(2)_{C}$, leaving the combination $U(1)_{A}=U(1)_{H}-U(1)_{C}$ as a distinguished global symmetry commuting with $U(1)_{R}$. This leads to the refined twisted half index,
@@ 2.1 Definition of the Macdonald index
Superconformal index counts certain BPS states with $(-1)^{F}$ or, equivalently, short multiplets modulo recombinations Kinney:2005ej . Because of the Bose-Fermi cancellation, this quantity is invariant under marginal deformations. The trace formula for the full superconformal index of 4d $\mathcal{N}=2$ SCFT is defined as
@@
where $t=qT$ and the trace is taken over the states satisfying $\Delta=2R+2j_{2}-r$ and $j_{1}-j_{2}+r=0$, and it counts $\frac{1}{4}$-BPS states.
@@
This is referred to as the Schur index Gadde:2011ik , which plays a pivotal role in many developments in the field. In particular, the Schur index computes the (super-)character of the associated VOA of Beem:2013sza . One can also directly obtain the Schur index by taking the $q=t$ limit from the full index. In this limit, the $p$-dependence drops out Gadde:2011uv . One interesting aspect of the Schur index is that, unlike the Macdonald index, the trace can be taken over all states in the radially quantized Hilbert space, without restricting to a certain BPS subspace.
@@ 2.2 Computation of Macdonald index
For a Lagrangian theory, the superconformal index can be computed simply by combining contributions from the free fields and projecting them to the gauge-invariant sector by integrating over the gauge group. However, for a non-Lagrangian theory, such as the Argyres-Douglas theory Argyres:1995jj ; Argyres:1995xn ; Eguchi:1996vu ; Cecotti:2010fi ; Xie:2012hs , we need to resort to alternative methods. The Macdonald index for Argyres-Douglas theories can be computed using several methods, each with its own strengths and weaknesses. Also, each methods give rise to a different formula, which is supposed to be identical in the end. This provides non-trivial mathematical identities.
@@
The AD theories of $(A_{1},A_{n})$, $(A_{1},D_{n})$, $(A_{k-1},A_{kn-1})$, $(A_{k-1},D_{kn+1})$ type and a few more admit $\mathcal{N}=1$ Lagrangian gauge theories flow to $\mathcal{N}=2$ fixed points Maruyoshi:2016tqk ; Maruyoshi:2016aim ; Agarwal:2016pjo ; Agarwal:2017roi ; Benvenuti:2017bpg . Therefore, one can use the contour integral formula for the full index and take the Macdonald limit to obtain the index. So far, this is the only known way to obtain the full index of AD theories.
However, this method has the following drawback: Firstly, not all AD theories admit such a Lagrangian description. For example, $(A_{k-1},A_{n-1})$ theory with $(k,n)=1$, whose associated VOA is given by the $W$-minimal model $W(k,k+n)$, does not have a known Lagrangian description for $k>1$. Secondly, this method is computationally inefficient for obtaining the Macdonald index. Especially, the integrand is singular as we take the $p\to 0$ limit. We need to evaluate the integral first and then take the Macdonald limit.
@@
Class $\mathcal{S}$ theoery refers to 4d $\mathcal{N}=2$ SCFTs obtained via compactifying 6d $\mathcal{N}=(2,0)$ theory on a Riemann surface Witten:1997sc ; Gaiotto:2009we ; Gaiotto:2009hg . Such a 4d theory is determined by the choice of 6d theory ($\Gamma\in ADE$) and a UV curve $\mathcal{C}_{g,n}$ of genus $g$ and $n$ punctures. In addition, one should prescribe a local structure for each puncture. Many of the AD theories can be realized in class $\mathcal{S}$ Xie:2012hs ; Wang:2015mra .
@@
One can employ SCFT/VOA correspondence to compute the Schur index from the vacuum character of the associated vertex operator algebra Beem:2013sza . Argyres-Douglas theories have relatively simple associated VOA such as Virasoro/W-minimal models or affine Lie algebra Buican:2015ina ; Cordova:2015nma ; Xie:2016evu ; Song:2017oew .
Since VOA captures the entire Schur sector, it was expected that VOA would somehow capture the Macdonald index, as it counts the same set of operators. Indeed, it was found that the ‘refined character’ of the vacuum module can capture the refined index Song:2016yfd . See also Agarwal:2018zqi ; Bonetti:2018fqz ; Beem:2019tfp ; Xie:2019zlb ; Agarwal:2021oyl . For the case of Virasoro algebra, we consider the following filtration of the vacuum module
@@
This formula agrees with (9) to high orders in $q$. The conjecture is that both formulae are identical so that the refined character defined as above indeed computes the Macdonald index.
A fermionic formula for the $W(3,7)$ minimal model is also written in Foda:2019guo , which is identical to the Macdonald index of $(A_{2},A_{3})=(A_{1},E_{6})$ theory:
@@
This method involves computing Gröbner basis for the coordinate ring of the arc space (or jet scheme), which gives us a computational bottleneck. However, this method is sometimes more suitable to obtain fermionic expression for the index, as is demonstrated in bai2020quadratic ; bruschek2011arc ; Andrews:2025krn ; Kang:2025zub .
@@ 2.3 Conjectural formula for the Macdonald index from IR formula
The main result of this paper is a conjectural formula for the Macdonald index, given as a certain modification of the IR formula of Cordova:2015nma obtained by evaluating the trace of an operator $\mathcal{O}$ (called the Kontsevich-Soibelman operator) constructed out of the BPS particle spectrum in the Coulomb branch. The BPS particle spectrum varies along the Coulomb branch moduli space, in particular as it crosses a certain codimension 1 wall subspace. A crucial property of the operator $\mathcal{O}$ is that it is wall-crossing invariant, so that it is an invariant of the 4d theory. It was found that the trace of $\mathcal{O}$ computes the Schud index of the 4d theory. Let us briefly review their formula.
@@
For each charge vector $\gamma$ of the 4d theory, we introduce a torus algebra variable $X_{\gamma}$ which obey
$X_{\gamma}X_{\gamma^{\prime}}=q^{\frac{1}{2}\left<\gamma,\gamma^{\prime}\right>}X_{\gamma+\gamma^{\prime}}=q^{\left<\gamma,\gamma^{\prime}\right>}X_{\gamma}X_{\gamma^{\prime}}.$ (20)
@@
$E_{q}(z)=(q^{\frac{1}{2}}z;q)_{\infty}^{-1}=\sum_{n\in\mathbb{Z}_{\geq 0}}\frac{(q^{\frac{1}{2}}z)^{n}}{(q;q)_{n}}$ (22)
@@
As shown in Cordova:2015nma , this IR formula indeed reproduces the Schur index. Motivated by the IR formula, we propose a Macdonald index formula that involves a certain tweak.
@@
Let us describe our conjectural formula for the Macdonald index of $(A_{1},A_{2n})$ Argyres-Douglas theory. To begin with, let us consider the IR formula for the Schur index.
In the canonical chamber, the Schur index is computed from the KS operator
$\mathcal{O}(q)$ $=\prod_{i\textrm{~odd}}E_{q}(X_{\gamma_{i}})\prod_{j\textrm{~even}}E_{q}(X_{\gamma_{j}})\prod_{i\textrm{~odd}}E_{q}(X_{-\gamma_{i}})\prod_{j\textrm{~even}}E_{q}(X_{-\gamma_{j}})\ ,$ (25)
upon multiplying the overall free vector piece $(q;q)^{2k}$. That is, we first put the source parts before the sink part in the BPS quiver. The trace of the KS operator can be evaluated to give
$\begin{split}\mathrm{Tr}\mathcal{O}(q)&=\sum_{\ell_{i},m_{i}}\frac{(q^{\frac{1}{2}})^{\ell_{1}+m_{1}+\cdots+\ell_{n}+m_{n}}}{\prod_{i=1}^{n}(q)_{\ell_{i}}(q)_{m_{i}}}\mathrm{Tr}X_{\gamma_{1}}^{\ell_{1}}\cdots X_{\gamma_{n}}^{\ell_{n}}X_{-\gamma_{1}}^{m_{1}}\cdots X_{-\gamma_{n}}^{m_{n}}\\ &=\sum_{\ell_{i},m_{i}}\frac{q^{\frac{\sum_{i=1}^{n}\ell_{i}+m_{i}}{2}+\sum_{i=1}^{n-1}\ell_{i}m_{i+1}}}{\prod_{i=1}^{n}(q)_{\ell_{i}}(q)_{m_{i}}}\mathrm{Tr}X_{\gamma_{1}}^{\ell_{1}}X_{-\gamma_{1}}^{m_{1}}\cdots X_{\gamma_{n}}^{\ell_{n}}X_{-\gamma_{n}}^{m_{n}}\ ,\end{split}$ (26)
where we used $X_{\gamma}X_{\gamma^{\prime}}=q^{\langle\gamma,\gamma^{\prime}\rangle}X_{\gamma^{\prime}}X_{\gamma}$. Upon imposing $\mathrm{Tr}X_{\gamma}^{\ell}X_{-\gamma}^{m}=\delta_{\ell,m}$, we reproduce the Schur index of the AD theory, given as
$I_{(A_{1},A_{2n})}(q)=(q)_{\infty}^{2n}\sum_{m_{1},\cdots,m_{2n}\geq 0}\frac{q^{\sum_{i=1}^{2n}m_{i}+\sum_{i=1}^{2n-1}m_{i}m_{i+1}}}{\prod_{i=1}^{2n}(q)^{2}_{m_{i}}}\ .$ (27)
It is proven in Cecotti:2015lab that this formula is indeed identical to the vacuum character of the Virasoro minimal model $M(2,2n+3)$.
From this expression, we propose the following conjectural formula for the Macdonald index:
$\boxed{I_{\textrm{Mac}}^{A_{2n}}(q,t)=(q)_{\infty}^{n}(t)_{\infty}^{n}\sum_{m_{1},\ldots,m_{2n}\geq 0}\frac{q^{\sum_{i=1}^{2n}m_{i}+\sum_{i=1}^{2n-1}m_{i}m_{i+1}}}{\prod_{i\textrm{~odd}}(q)_{m_{i}}^{2}\prod_{j\textrm{~even}}(q)_{m_{j}}(t)_{m_{j}}}T^{\sum_{i=1}^{n}m_{2i-1}}}$ (28)
where $t=qT$. Here, we replaced the prefactor that comes from $n$ copies of abelian vector multiplets $(q)_{\infty}^{2n}$ by that of the Macdonald index $(q)_{\infty}^{n}(t)_{\infty}^{n}$, and introduced the Macdonald grading fugacity $T$ for the ‘sink nodes’ ($n_{1},n_{3},\cdots$). In addition, we replaced the $(q)_{m}^{2}$ in the demoninator by $(q)_{m}(t)_{m}$ for the ‘source nodes’ ($n_{2},n_{4},\cdots$).
It has been suspected that there should exist an IR formula analogous to Cordova:2015nma for the Macdonald index since the Schur and Macdonald indices count the same set of states. This is also motivated by the observation that $E_{q}(X)$ is identical to the Schur index of a free hypermultiplet. Therefore, one can think of replacing $E_{q}(X)$ by $E_{q,t}(X)$ given as Song:2016yfd
which is the same as the Macdonald index for a ‘free’ hypermultiplet. By observation, we find that at least in the canonical chamber, where all the nodes are either sink or source, we find the following ad hoc IR formula:
[…]
That is, we have separated source/sink nodes to be distinct. We replaced $E_{q}(X_{\gamma})$ by $E_{q,t}(X_{\gamma})$ for the odd (source) nodes, and replaced $E_{q}(X_{\gamma})$ by $E_{q}(\widehat{X}_{\gamma})$ for the even (sink) nodes. To evaluate the index, we first need to expand the above expression in terms of $q$-series and collect all the generators $X_{\gamma}$’s and $\widehat{X}_{\gamma}$’s to the right. Namely, we write in the form of
$\mathcal{O}(q,t)=\sum_{\vec{n}}f(q,t,n)\,\mathrm{Tr}\left(\prod_{i}X_{\gamma_{i}}^{n_{i}}\prod_{j}\widehat{X}_{\gamma_{j}}^{n_{j}}\right)\ .$ (31)
Once we arrive at this form, we impose the normalization for the trace of $\widehat{X}_{\gamma}$ to be
$\mathrm{Tr}\widehat{X}_{\gamma}^{\ell}\widehat{X}_{-\gamma}^{m}=\delta_{\ell,m}\frac{(q;q)_{\ell}}{(t;q)_{\ell}}\ .$ (32)
We experimentally find that this ‘IR formula’ works for a certain fixed chamber in a number of theories. Let us emphasize that our prescription is well-defined only in a particular chamber. We do not have a direct justification of this formula, but we will give a physical argument for the formula (28) using the 3d twisted reduction of 4d theory.
$(A_{1},A_{2n+1})$ theory
We propose the Macdonald index for the $(A_{1},A_{2n+1})$ Argyres-Douglas theory following the same prescription we described above as
@@
Following the same procedure, we conjecture the Macdonald index for the $(A_{1},D_{2n+1})$ Argyres-Douglas theory is given by the following expression:
@@
Here $z$ is the $SU(2)$ flavor fugacity. We find this expression to agree with the Macdonald index computed using other methods in high orders in $q$.
$(A_{1},D_{2n+2})$ theory
Similarly, we find the following expression for the Macdonald index of $(A_{1},D_{2n+2})$ theory:
Here $z_{1},z_{2}$ are the flavor fugacities of $SU(2)\times U(1)$ flavor symmetry. Here, $z_{1}$ and $z_{2}$ correspond to some linear combinations of the Cartans. The precise map will be given in Section 4.4. We have checked this expression to agree with known ones to high orders in $q$.
@@
3 Methods
In this section, we describe our strategy to obtain a fermionic sum formula for the Macdonald index via twisted dimensional reduction to a 3d theory. The procedure is as follows:

### v2

@@ Abstract
We propose a new fermionic sum formula for the Macdonald index of a class of Argyres-Douglas theories. The formula arises naturally from a three-dimensional topological field theory obtained via a twisted dimensional reduction of the 4d theory. Such a reduction often gives rise to a 3d $\mathcal{N}=2$ abelian Chern-Simons matter theory, which is expected to flow to an $\mathcal{N}=4$ superconformal fixed point. After performing a topological twist, we obtain a 3d TFT admitting boundary conditions that support the vertex operator algebra associated with the original 4d theory. In this framework, the Macdonald index appears as a half-index of the 3d gauge theory, with the Macdonald grading determined by a distinguished $U(1)_{A}$ symmetry in the infrared $\mathcal{N}=4$ superconformal algebra. We present a general procedure to identify this $U(1)_{A}$ symmetry and, whenever possible, show that it reproduces the refined character of the associated vertex operator algebra, thereby recovering the Macdonald index. Our construction also gives a hint towards the IR formula for the Macdonald index in terms of 4d BPS particles on the Coulomb branch.
@@ 1 Introduction
$I_{S}(q)=(q)_{\infty}^{2r}\text{Tr }{\mathcal{O}}(q)\ ,$ (2)
where $r$ is the rank of the Coulomb branch and ${\mathcal{O}}(q)$ is the quantum monodromy operator constructed from the BPS spectrum in the Coulomb branch of the 4d theory. This representation provides a deep connection between UV operator algebras and infrared BPS dynamics, suggesting interrelations among wall-crossing phenomena, protected operator algebras, and the underlying modular structure of supersymmetric field theories.
@@
where the trace is taken over the space of boundary local operators, and $R_{H},~R_{C}$ correspond to Cartans of the $SU(2)_{H}\times SU(2)_{C}$ R-symmetry of the $\mathcal{N}=4$ algebra. For a large class of Argyres-Douglas (AD) theories, the $A$-twisted half-index obtained in this way is identified with the trace of the quantum monodromy operator (2), which in turn provides a physical motivation for the infrared trace formula of the Schur index. See also ArabiArdehali:2024ysy ; Ardehali:2024 for an alternative approach based on studying the high-temperature limit of the superconformal index.
@@
where the trace is over the operators satisfying the relation $\Delta=2R+2j_{2}-r$ and $j_{1}-j_{2}+r=0$. This is a natural refinement to consider in the context of the 4d SCFT/2d VOA correspondence, since it is built from the same protected operator spectrum as the Schur index while keeping track of an additional $U(1)$ charge. From the perspective of the associated VOA, this refinement arises from the existence of a natural filtration structure, as first discussed in Song:2016yfd . See also Agarwal:2018zqi ; Bonetti:2018fqz ; Beem:2019tfp ; Xie:2019zlb ; Agarwal:2021oyl for related development.
The 3d $\mathcal{N}=4$ SCFT obtained from the twisted circle reduction has an $\mathcal{N}=2$ superconformal subalgebra where the $U(1)_{R}$ symmetry embeds into $SU(2)_{H}\times SU(2)_{C}$, leaving the combination $U(1)_{A}=U(1)_{H}-U(1)_{C}$ as a distinguished global symmetry commuting with the $U(1)_{R}$. This leads to the refined twisted half-index,
@@ 2.1 Definition of the Macdonald index
The superconformal index counts certain BPS states with weight of $(-1)^{F}$ or, equivalently, short multiplets modulo recombinations Kinney:2005ej . Because of the Bose-Fermi cancellation, this quantity is invariant under marginal deformations. The trace formula for the full superconformal index of a 4d $\mathcal{N}=2$ SCFT is defined as
@@
where $t=qT$ and the trace is taken over states satisfying $\Delta=2R+2j_{2}-r$ and $j_{1}-j_{2}+r=0$, counting $\frac{1}{4}$-BPS states.
@@
This is referred to as the Schur index Gadde:2011ik , which plays a pivotal role in many developments in the field. In particular, the Schur index computes the (super-)character of the associated VOA, in the context of 4d SCFT/VOA correspondence Beem:2013sza . One can also directly obtain the Schur index by taking the $q=t$ limit from the full index. In this limit, the $p$-dependence drops out Gadde:2011uv . One interesting aspect of the Schur index is that, unlike the Macdonald index, the trace can be taken over all states in the radially quantized Hilbert space, without restricting to a certain BPS subspace.
@@ 2.2 Computation of Macdonald index
For a Lagrangian theory, the superconformal index can be computed simply by combining contributions from the free fields and projecting them to the gauge-invariant sector by integrating over the gauge group. However, for a non-Lagrangian theory, such as the Argyres-Douglas theory Argyres:1995jj ; Argyres:1995xn ; Eguchi:1996vu ; Cecotti:2010fi ; Xie:2012hs , we need to resort to alternative methods. The Macdonald index for Argyres-Douglas theories can be computed using several methods, each with its own strengths and weaknesses. Moreover, each method yields a different formula, though the final expressions are expected to coincide, leading to non-trivial mathematical identities.
@@
The AD theories of $(A_{1},A_{n})$, $(A_{1},D_{n})$, $(A_{k-1},A_{kn-1})$, $(A_{k-1},D_{kn+1})$ type and a few more admit $\mathcal{N}=1$ Lagrangian gauge theory descriptions that flow to $\mathcal{N}=2$ fixed points Maruyoshi:2016tqk ; Maruyoshi:2016aim ; Agarwal:2016pjo ; Agarwal:2017roi ; Benvenuti:2017bpg . Therefore, one can use the contour integral formula for the full index and take the Macdonald limit to obtain the index. So far, this is the only known way to obtain the full index of AD theories.
However, this method has some drawbacks: Firstly, not all AD theories admit such a Lagrangian description. For example, $(A_{k-1},A_{n-1})$ theory with $(k,n)=1$, whose associated VOA is given by the $W$-minimal model $W(k,k+n)$, does not have a known Lagrangian description for $k>2$. Secondly, this method is computationally inefficient for obtaining the Macdonald index. In particular, the integrand is singular as we take the $p\to 0$ limit. So one must first evaluate the integral of the full index and then take the Macdonald limit.
@@
Class $\mathcal{S}$ theory refers to 4d $\mathcal{N}=2$ SCFTs obtained via compactifying 6d $\mathcal{N}=(2,0)$ theory on a Riemann surface Witten:1997sc ; Gaiotto:2009we ; Gaiotto:2009hg . Such a 4d theory is determined by the choice of 6d theory ($\Gamma\in ADE$) and a UV curve $\mathcal{C}_{g,n}$ of genus $g$ and $n$ punctures. In addition, one should prescribe a local structure for each puncture. Many of the AD theories can be realized in class $\mathcal{S}$ Xie:2012hs ; Wang:2015mra .
@@
One can employ the SCFT/VOA correspondence to compute the Schur index from the vacuum character of the associated vertex operator algebra Beem:2013sza . Argyres-Douglas theories have relatively simple associated VOAs such as Virasoro/W-minimal models or affine Lie algebra Buican:2015ina ; Cordova:2015nma ; Xie:2016evu ; Song:2017oew .
Since a VOA captures the entire Schur sector, it was expected that the VOA would somehow capture the Macdonald index, which counts the same set of operators. Indeed, it was found that the ‘refined character’ of the vacuum module can capture the refined index Song:2016yfd . See also Agarwal:2018zqi ; Bonetti:2018fqz ; Beem:2019tfp ; Xie:2019zlb ; Agarwal:2021oyl . For the case of Virasoro algebra, we consider the following filtration of the vacuum module
@@
One can check that this formula agrees with (9) to high orders in $q$. The conjecture is that they are are identical, confirming that the refined character defined as above indeed computes the Macdonald index.
A fermionic formula for the $W(3,7)$ minimal model is also proposed in Foda:2019guo , which computes the Macdonald index of $(A_{2},A_{3})=(A_{1},E_{6})$ theory:
@@
This method involves computing a Gröbner basis for the coordinate ring of the arc space (or jet scheme), which gives us a computational bottleneck. However, this method is sometimes more suitable for obtaining fermionic expressions for the index, as is demonstrated in bai2020quadratic ; bruschek2011arc ; Andrews:2025krn ; Kang:2025zub .
@@ 2.3 Conjectural formula for the Macdonald index from IR formula
It is conjectured in Cordova:2015nma that the Schur index can be computed by evaluating the trace of an operator $\mathcal{O}$ in quantum torus algebra, known as the Kontsevich-Soibelman (KS) operator, constructed out of the BPS particle spectrum in the Coulomb branch of the 4d theory. The spectrum of BPS particles depends on the point on the Coulomb branch and can jump across codimension-one walls of marginal stability. A key property of the operator $\mathcal{O}$ is that it remains invariant under wall-crossing, and therefore it defines an invariant of the 4d theory.
@@
Let $\Gamma$ be the electro-magnetic charge lattice of the 4d Coulomb branch. For each charge vector $\gamma\in\Gamma$, we introduce a quantum torus algebra variable $X_{\gamma}$ which obeys
$X_{\gamma}X_{\gamma^{\prime}}=q^{\frac{1}{2}\left<\gamma,\gamma^{\prime}\right>}X_{\gamma+\gamma^{\prime}}=q^{\left<\gamma,\gamma^{\prime}\right>}X_{\gamma^{\prime}}X_{\gamma}.$ (20)
@@
$E_{q}(X)=(q^{\frac{1}{2}}X;q)_{\infty}^{-1}=\sum_{n\in\mathbb{Z}_{\geq 0}}\frac{(q^{\frac{1}{2}}X)^{n}}{(q;q)_{n}}$ (22)
@@
As shown in Cordova:2015nma , this IR formula indeed reproduces the Schur index.
For the Argyres-Douglas theories of type $(A_{1},G)$, there exists the so-called canonical chamber, where all the nodes are either sink or source. In this chamber the formula reads
$\mathcal{O}(q)$ $=\prod_{i\textrm{~odd}}E_{q}(X_{\gamma_{i}})\prod_{j\textrm{~even}}E_{q}(X_{\gamma_{j}})\prod_{i\textrm{~odd}}E_{q}(X_{-\gamma_{i}})\prod_{j\textrm{~even}}E_{q}(X_{-\gamma_{j}})\ .$ (25)
For example, for $(A_{1},A_{2n})$ theory, we obtain a fermionic sum type expression:
$I_{\text{Schur}}(q)=(q)_{\infty}^{2n}\sum_{m_{1},\cdots,m_{2n}\geq 0}\frac{q^{\sum_{i=1}^{2n}m_{i}+\sum_{i=1}^{2n-1}m_{i}m_{i+1}}}{\prod_{i=1}^{2n}(q)^{2}_{m_{i}}}\ .$ (26)
It is proven in Cecotti:2015lab that this formula agrees with the vacuum character of the Virasoro minimal model $M(2,2n+3)$.
It has been suspected that an analogous IR formula should exist for the Macdonald refinement. This is partly motivated by the observation that $E_{q}(X)$ is identical to the Schur index of a free hypermultiplet. Therefore, one can consider replacing $E_{q}(X)$ by $E_{q,t}(X)$ given as Song:2016yfd
which indeed computes the Macdonald index of the free hypermultiplet theory. Empirically, we find that the following ad hoc modification computes the Macdonald refinement:
[…]
where $(t)_{n}$ denotes $(t;q)_{n}$. Here we replaced $E_{q}(X_{\gamma})$ by $E_{q,t}(X_{\gamma})$ for the sink nodes, and replaced $E_{q}(X_{\gamma})$ by $\widehat{E}_{q}({X}_{\gamma})$ for the anti-particles in the source nodes, where
$\widehat{E}_{q}(X)=\sum_{m\geq 0}\frac{(q^{\frac{1}{2}}X)^{m}}{(t;q)_{m}}\ .$ (29)
We experimentally find that this ‘refined IR formula’ applies to a large class of $(A_{1},G)$ type theories in the canonical chamber. Let us emphasize that this prescription is well-defined only in this particular chamber. While we do not have a direct justification for this formula, we will give a physical argument for the formula based on the three-dimensional twisted reduction of 4d theory in the following sections.
@@
For the $(A_{1},A_{2n})$ Argyres-Douglas theory, the proposed refinement procedure gives the following conjectural formula for the Macdonald index:
$\boxed{I_{\textrm{Mac}}^{A_{2n}}(q,t)=(q)_{\infty}^{n}(t)_{\infty}^{n}\sum_{m_{1},\ldots,m_{2n}\geq 0}\frac{q^{\sum_{i=1}^{2n}m_{i}+\sum_{i=1}^{2n-1}m_{i}m_{i+1}}}{\prod_{i\textrm{~even}}(q)_{m_{i}}^{2}\prod_{j\textrm{~odd}}(q)_{m_{j}}(t)_{m_{j}}}T^{\sum_{i=1}^{n}m_{2i}}\ ,}$ (30)
where $t=qT$. We find that this formula agrees with the Macdonald index computed via other methods to very high orders in $q$. If we compare this with the Schur index, the prefactor that comes from $n$ copies of abelian vector multiplets $(q)_{\infty}^{2n}$ is replaced by that of the Macdonald index $(q)_{\infty}^{n}(t)_{\infty}^{n}$, and the Macdonald grading fugacity $T$ is introduced for the ‘sink nodes’ ($n_{2},n_{4},\cdots$). In addition, the factor of $(q)_{m}^{2}$ in the denominator is replaced by $(q)_{m}(t)_{m}$ for the ‘source nodes’ ($n_{1},n_{3},\cdots$).
@@
Following the same procedure, we conjecture that the Macdonald index for the $(A_{1},D_{2n+1})$ Argyres-Douglas theory is given by the following expression:
@@
Here $z$ is the $SU(2)$ flavor fugacity. We find that this expression agrees with the Macdonald index computed via other methods to high orders in $q$.
@@
$(A_{1},A_{2n+1})$ theory
We propose the Macdonald index for the $(A_{1},A_{2n+1})$ Argyres-Douglas theory following the same prescription, but here the roles of the sinks and sources are interchanged
[…]
This gives the following expression for the Macdonald index
$(A_{1},D_{2n+2})$ theory
Similarly, with (35), we find the following expression for the Macdonald index of $(A_{1},D_{2n+2})$ theory:
Here $z_{1},z_{2}$ are the flavor fugacities of $SU(2)\times U(1)$ flavor symmetry. Here, $z_{1}$ and $z_{2}$ correspond to some linear combinations of the Cartans. The precise map will be given in Section 4.4. We have checked that this expression agrees with known ones to high orders in $q$.
3 Macdonald index from the twisted compactification
In this section, we describe our strategy to justify the fermionic sum formula for the Macdonald index proposed in the previous section via twisted dimensional reduction to a 3d theory. The procedure is as follows: